OpenStudy (anonymous):

7 + 2y = 8x 3x - 2y = 0 Solve the system of equations by substitution.

OpenStudy (vishweshshrimali5):

find the value of y in terms of x from second eq. and put it in first

OpenStudy (vishweshshrimali5):

3x - 2y = 0 \[\large x = \cfrac{2y}{3}\] put this value of x in first eq. to get value of y

OpenStudy (anonymous):

First I would put one equation in in y=mx+b format y=4x-7/2 Use that and plug it into your other equation. 3x-2(4x-7/2)=0 From there work it out to find your x Once you've found your X, plug your x into original problem to solve for Y

OpenStudy (anonymous):

thanks a lot :)

OpenStudy (anonymous):

No problem (: